3x^2 - 4x - 7 = 0
and
x2 + 5x - 6 = 0
Very confused.

Found 2 solutions by checkley75, scott8148:
Answer by checkley75(3666)

3X^2-4X-7=0
(3X-7)(X+1)=0
3X-7=0
3X=7
X=7/3 ANSWER.
X+1=0
X=-1 ANSWER.
-----------------------------------------
X^2+5X-6=0
(X+6)(X-1)=0
X+6=0
X=-6 ANSWER.
X-1=0
X=1 ANSWER.

Answer by scott8148(6628)

factoring:

(3x-7)(x+1)=0

(x+6)(x-1)=0